OpenStudy (anonymous):

Okay, so I am currently learning about normal distribution, and I don't know how to do the question that looks like this (I'll try my best to convey it, but I don't have all the right symbols on my laptop): P(X is less than or equal to 64) The mean is 76 and the standard deviation is 6. - - Please no vague or unsure answers.

jimthompson5910 (jim_thompson5910):

x = 64 is the given raw score the first step is to convert this to a standard z-score z = (x - mu)/sigma z = (x - 76)/6 z = (64 - 76)/6 z = -12/6 z = -2 so the raw score of x = 64 corresponds to the standard z-score of z = -2 z = -2 means that the raw score x = 64 is 2 standard deviations below the mean (mu = 76)

jimthompson5910 (jim_thompson5910):

now use a table to compute P(z < -2) a table like this is usually found in the back of your stats book http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf

OpenStudy (anonymous):

THANK YOU SO VERY MUCH!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

you can also use the `normalcdf` function found on TI83, TI84, etc calculators http://tibasicdev.wikidot.com/normalcdf but your teacher will probably want you to use the table instead

OpenStudy (anonymous):

How about if you are finding p(x is GREATER THAN or equal to 70) - -same mean and standard deviation :)

jimthompson5910 (jim_thompson5910):

Rule: P(x > k) = 1 - P(x < k) where k is some fixed number

jimthompson5910 (jim_thompson5910):

using the rule, we know that P(x > 70) = 1 - P(x < 70)

jimthompson5910 (jim_thompson5910):

so you need to follow the steps to compute P(x < 70) first then subtract that result from 1

jimthompson5910 (jim_thompson5910):

oh and it doesn't matter if you have `greater than` or `greater than or equal to` since the two lead to the same answer

OpenStudy (anonymous):

Okay last question: what if it has p(x is greater than or equal to 85 but less than or equal to 91. - -I got .93 and .43 as the "Z's" for both of them, then I tried to subtract them from each other: .93-.43, but I think I did it wrong. (Sorry if that doesn't make any sense and also thank you so much for your help).

jimthompson5910 (jim_thompson5910):

Rule \[\Large P\left(a < x < b\right) = P(x < b) - P(x < a)\] in this case, `b > a`

jimthompson5910 (jim_thompson5910):

\[\Large P\left(85 \le x \le 91\right) = P(x \le 91) - P(x \le 85)\]

jimthompson5910 (jim_thompson5910):

Step 1) Compute \(\Large P(x \le 91)\). Call this result M Step 2) Compute \(\Large P(x \le 85)\). Call this result N Step 3) Compute `M - N`